The kinetic energy of rotation of a
particle in circular motion is 100 J. It moves
once in the circle in 44s. Calculate the
rotational inertia of the particle.
1
Expert's answer
2018-06-24T10:12:08-0400
The kinetic energy of rotation of a particle:
E=1/2 Iω^2
ω=2π/44 rad/s.
I=2E/ω^2 =2(100)/(2π/44)^2 = 9800 kgm^2